OpenStudy (anonymous):

i need this ASAP: assume that the sales of a certain appliance dealer are approximated by a linear function.. suppose that sales were $13,500, in 1982 and $52,500 in 1987. let x=0 represent 1982. find the equation giving yearly sales S(x).

OpenStudy (anonymous):

@thushananth01

OpenStudy (anonymous):

anyone?? :\

OpenStudy (campbell_st):

find the slope of the line between (0, 13500) and (5, 52500) the y intercept occurs when x = 0 so you'll be able to put the line in the form y = mx + b

OpenStudy (anonymous):

\[S(x)=7800~~?\]

OpenStudy (campbell_st):

ummm nope thats no my answer... what slope did you get

OpenStudy (campbell_st):

thats the slope... but whats the equation...

OpenStudy (anonymous):

\[{52500-13500\over5-0}~~~~~=7800\]

OpenStudy (anonymous):

y=7800x+0

OpenStudy (anonymous):

y(0) =13500

OpenStudy (anonymous):

y=S(x) and b=0 S(x)=7800x ?

OpenStudy (campbell_st):

nope you need the y-intercept....

OpenStudy (anonymous):

when x = 0 y= 13500

OpenStudy (campbell_st):

but thats a good start

OpenStudy (anonymous):

52500=7800(5)+b 52500=3900+b b=48600 y=7800x+48600

OpenStudy (campbell_st):

ummm I think you're missing a zero on the right hand side...

OpenStudy (campbell_st):

the equation is S(x) = 7800x + 13500

OpenStudy (anonymous):

oh right sorry its S(x)=7800x+13500

OpenStudy (campbell_st):

b is the y intercept... and occurs when x = 0

OpenStudy (campbell_st):

good job

OpenStudy (anonymous):

thanks...i have another question...help please